Circular No. 6780

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SATELLITES OF URANUS Further selected CCD observations are given below, for S/1997 U 1

and S/1997 U 2, respectively. As on IAUC 6764 and 6771, the selection is restricted to one observation per satellite per observer per night.

1997 UT  α_{2000}  δ_{2000}  Observer
Nov. 26.20765 20^{h}32^{m}16.^{s}03 −19^{o}33′16′′.3 Lilly
27.20229 20 32 24.32 −19 32 46.9  _{”}
28.20530 20 32 32.82 −19 32 16.5  _{”}
Nov. 10.09307 20^{h}30^{m}07.^{s}18 −19^{o}32′00′′.9 Hergenrother
25.08604 20 31 55.75 −19 25 16.7 Offutt

S. Lilly, M.-C. Hainaut, O. Le Fèvre and M. Dussert (Mauna Kea). Canada-

France-Hawaii Telescope. Measurer B. Gladman. C. W. Hergenrother (Mt. Hopkins). 1.2-m f/8 reflector. W. Offutt (Cloudcroft). 0.6-m f/7 Ritchey-Chrétien.

Further computations by B. G. Marsden and G. V. Williams, Harvard-Smithsonian Center for Astrophysics, seem—surprisingly, in the case of the first, fainter satellite—to confirm the general correctness of the orbit solutions given on IAUC 6765. It appears that both objects really are Uranian satellites, and that their orbits (referred to the ecliptic) are indeed retrograde. It also appears that the orbit of the fainter satellite is significantly less eccentric than that of the brighter. The particular uranicentric solutions given here, for S/1997 U 1 and S/1997 U 2, respectively, have assumed eccentricities, deliberately chosen near the maximum possible value for the first satellite and near the minimum possible value for the second. The satellites should now be recoverable reasonably close to prediction on their return to the morning sky after conjunction.

T = 1997 June 27.506 TT  ω = 141.^{o}851
e = 0.20000  Ω = 185.758 }2000.0
q = 0.04158 AU  i = 146.360
a = 0.05197 AU n^{o} = 0.55008 P = 654 days
T = 1997 Feb. 10.299 TT  ω = 256.^{o}248
e = 0.40000  Ω = 221.244 }2000.0
q = 0.02590 AU  i = 153.362
a = 0.04317 AU n^{o} = 0.72657 P = 495 days
1997 November 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden